Operations Report – May 6th

Crew 235 Operations Report 6-MAY-2021

SOL: 4

Name of person filing report: Atila Meszaros

Non-nominal systems: ATV-2

Notes on non-nominal systems: ATV-2 gas pedal system is broken. Once you press it, it would get stuck, instead of retracting, and it could lead to an accident. There’s also a shift problem, but we suspect is related to the gas pedal system interfering with the shift change. As we leave on Saturday, the ATVs will have to be tuned up before next season.

Generator: Working nominally.

Hours run: 0 hours

From what time last night: the generator didn’t run

To what time this morning: the generator didn’t run

List any additional daytime hours when the generator was run: N/A

Solar— SOC 92% at 20:09 (Before generator is run at night)

Notes on power system: Nothing to report

Diesel Reading – 47%

Station Propane Reading – 80%

Outpost Propane Reading – 1. 70% 2. 80%

Water (loft tank): 25 gallons

Water Meter: 151161.3 units

Water (static tank): 250 gallons

Static to Loft Pump used – yes

Water in GreenHab: 0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: No

Sojourner rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Spirit rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Opportunity rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Curiosity rover used: No, in town

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Perseverance rover used: Yes

Hours: 211.9

Beginning charge: 100%

Ending charge: 31%

Currently charging: Yes

Notes on rovers: None

ATV’s Used: Yes

Reason for use: EVA#4 and EVA#5

Oil Added? No

ATV Fuel Used: N/A

# Hours the ATVs were used today: 2 hours

Notes on ATVs: ATV 2 gs pedal out of service

HabCar used and why, where?: Yes, to go to town for supplies.

CrewCar used and why, where?: Nothing to report

General notes and comments: Nothing to report

Summary of internet: Working nominally.

Summary of suits and radios: All radios nominal, all functional suits currently charging.

Summary of Hab operations: Mouse found in the morning and released during EVA #5.

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: General cleaning and bacterial disposal.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Battery replaced and now rotation system is working.

Summary of health and safety issues: Nothing to report

Questions, concerns, and requests to Mission Support: Thank you for being with us tonight, Bernard!

EVA Report – May 6th

Crew 235 EVA Report 06-May-2021

EVA # 6, 7

Author: Jen Carver-Hunter

Purpose of EVA: Continuation of rock sample collection

EVA 6 Start time: 0800
EVA 6 End time: 0900
EVA 7 Start Time: 0930
EVA 7 End time: 1030

Narrative: One of our crew’s goals is to collect rock samples that replicate the rocks found on Mars. For today’s EVA, our targeted specimens were gypsum and sandstone. The first planned stop for the EVA was the overlook at Copernicus Hwy, and the second planned stop on return to the Hab was Cowboy Corner.

We also decided to use these EVAs as an opportunity for the new crew to conduct an EVA without an MDRS-experienced leader. For this reason, we included a map lesson refresher prior to the EVA to ensure that all crew members would be familiar with the geologic features.

The first EVA crew (EVA 6) left the Hab in vehicles at 0800 and followed Cow Dung Rd north to Brahe Hwy. They missed the turn to Brahe Hwy and found themselves at the quarry, where they turned around and made their way to Brahe Hwy.

Then they followed Brahe Hwy to Copernicus Hwy where they stopped for a chance for a 360-degree look at the formations in the area. Then they reboarded vehicles and returned east on Brahe Hwy toward Cow Dung Road until they spotted the glint of the sun shining off the gypsum on the ground.

They collected samples of gypsum near Beige Moon before continuing on to Cowboy Corner, where they decided to not stop for sandstone collection. They also needed to release a packrat near the reservoir. After releasing it, they returned to vehicles and came back to the Hab without incident.

The second EVA crew (EVA 7) left the Hab in vehicles at 0930 and also followed Cow Dung Road north to Brahe Hwy. They did not miss the turn, so they executed a perfect left turn toward Copernicus Hwy, and stopped at the intersection of the two roads to take in the sights at the overlook.

The crew then started up their vehicles (one rover and one ATV) to start their trip back to the Hab. They also stopped near Beige Moon to collect gypsum samples. They elected to also stop at Cowboy Corner to collect sandstone samples before returning to the Hab without incident.

Destination: The overlook at Copernicus Hwy/Brahe Hwy; Beige Moon, Cowboy Corner

Coordinates (use UTM NAD27 CONUS): Destination 1 – 516100, 4254400; Destination 2 – 516500, 4254400; Destination 3 – 518350, 4252700

EVA 4 Participants: Jeff, Randall, Thomas

EVA 5 Participants: Krysta, Jennifer, Allison

Road(s) and routes per MDRS Map: Entrance Rd to Cow Dung Rd to Brahe Hwy

Mode of travel: Pathfinder Rover and 2 ATVs (only 1 ATV used on the second EVA)

Journalist Report – May 6th

Crew 235 Journalist Report 06May2021
Author: Jeff Streba, Crew Journalist
Sol 4: Log Entry

Another beautiful day on Mars.
We had an early morning EVA out to the overlook and then we collected Gypsum and Sandstone samples. Because of the lack of functional suits, only 3 people went on
each EVA.

While one EVA team was out the other team analyzed bacteria. We were all impressed with the results.

When everyone was together, we discussed how to incorporate both the bacteria experiment and rockhounding into our classrooms.

We then ended our sim

Copyright © The Mars Society. All rights reserved. | Main Site